Can you show me the steps to solve this please..
x-y-z=6
x-2y-3z=14
x-3y-7z=28
All of these three equation are used to find out the value of x, y, and z
Thank you very much..

Answer by stanbon(75887)

x-y-z=6
x-2y-3z=14
x-3y-7z=28
-------------------
Subtract the 1st Eq. from the 2nd
Subtract the 1st Eq. from the 3rd.
---
x-y-z=6
0-y-2z=8
0-2y-6z=22
-------------------------
Subtract 2 times the 2nd Eq from the 3rd.
x-y-z=6
0-y-2z=8
0-0-2z=6
---------------------------------
Solve the 3rd equation for "z":
-2z = 6
z = -3
------------------------------
Substitute z = -3 into the 2nd Eq. and solve for "y":
-y-2(-3) = 8
-y + 6 = 8
y = -2
----------------------
Substitute z=-3 and y=-2 into the 1st Eq and solve for "x":
x-(-2)-(-3) = 6
x+2+3 = 6
x = 1
---------------------------------
Cheers,
Stan H.